OpenStudy (anonymous):

The Formula for the perimiter of a rectangle is p=2l+2w, where l=Length and W=width. Cindy has 14 meters of fencing. SHe plans to Fence in a rectangular dog run that is 2 meters wide. How long can she make the run??

OpenStudy (anonymous):

P = 2l + 2w 14 = 2l + 2(2) 14 = 2l + 4 10 = 2l Can you finish solving for l?

OpenStudy (anonymous):

@Calcmathlete is it 12?? or 5??

OpenStudy (anonymous):

5

OpenStudy (anonymous):

ok i see....thankss!